OpenStudy (anonymous):

The endpoints of one diagonal of a rhombus are (0, -8) and (8, -4). If the coordinates of the 3rd vertex are (1, 0), what are the coordinates of the 4th vertex?

OpenStudy (anonymous):

sketch graph from y = -8 must go down 4 so at y = -12 from x = 0 must go right 7 so at x = 7 so (7,-12)

OpenStudy (anonymous):

Credit goes to damon: http://www.jiskha.com/display.cgi?id=1326927624

OpenStudy (amistre64):

http://openstudy.com/users/jewjewbri#/updates/4fc944d7e4b0c6963ad3a374 gotta start reading a new book :/

OpenStudy (anonymous):

First, find the slope of the diagonal which is 1/2. Then find the slope of a segment that is perpendicular to this one which is -2. Then find the midpoint of the diagonal which is (4, -6). Then find the distance between the midpoint and the third vertex which is 3√5. Then plug that into the distance formula. 3√5 = √((x - 4)^2 + (y + 6)^2) 45 = ((x - 4)^2 + (y + 6)^2) 45 = x^2 - 8x + 16 + y^2 + 12y + 36 45 = x^2 - 8x + y^2 + 12y + 52 Then plug it back into the slope formula. -2 = (-6 - y)/(4 - x) -8 + 2x = -6 - y y = 2 - 2x Then plug this into the distance formula. 45 = x^2 - 8x + y^2 + 12y + 52 0 = x^2 - 8x + y^2 + 12y + 7 0 = x^2 - 8x + (2 - 2x)^2 + 12(2 - 2x) + 7 0 = x^2 - 8x + 4x^2 - 8x + 4 + 24 - 24x + 7 0 = 5x^2 - 40x + 35 0 = (5x - 5)(x - 7) x = 1 or 7 which is 7 using logic. Plug it back in. y = 2 - 2(7) y = -12 (7, -12) Convoluted way for the win!!!! I guess you could just graph it. :)

OpenStudy (amistre64):

whatever gets you thru the night :)

OpenStudy (anonymous):

lol ^_^